# Radio Map Prediction from Aerial Images and Application to Coverage Optimization

### v1

@@ III-B Radio Map Simulations
For the Tx, we have selected the directional antenna type in the ray-tracing software. It allows to simulate an idealized main beam without any side lobes. By adjusting the parameters half power beam width and first null beam width, it can model very narrow beams or antennas covering wider sectors. For each Tx position, we test several combinations of azimuth and tilt angles in the direction pointing away from the building, on which the Tx is placed, with a line of sight algorithm. Orientations that do not cover at least a certain area on the ground are disregarded. For all angle combinations found, we select one of the narrow and one of the wide antenna patterns in Table II for separate simulations.
@@
The simulation output are the magnitudes of the channel coefficients at each Rx on the grid, which we use to calculate the path loss (1). The path loss values are converted to $\operatorname{dB}$-scale and rescaled and cut following [radiounet]. Assuming a bandwidth of $B=10\operatorname{MHz}$, thermal noise power spectral density $(N_{0})_{\operatorname{dBm}/\operatorname{Hz}}=-174\operatorname{dBm}/\operatorname{Hz}$ and an idealistic noise figure of $(\text{NF})_{\operatorname{dB}}=0\operatorname{dB}$ at the Rx (Table III), the noise floor $\mathcal{N}$ is calculated as
@@ III-C Other Data
We also include images from [geoportal_bb] taken about $2$ months after the LiDAR measurements, cut and downscaled to match the position and resolution of the nDSMs. Furthermore, the dataset contains 2D polygons with height attributes extracted from the nDSMs, which have been used for the ray-tracing simulations. Files containing line-of-sight information for each radio map are also included but not considered in this work.
IV EXPERIMENT DESIGN
IV-A CNN-Architectures
@@
As a lightweight baseline model, we use the RadioUNet [radiounet], more concretely the first part of the WNet described by the authors. In principle, it follows the structure of the original UNet [radiounet], but it features more down and upsampling layers and, in some parts, convolutions with a larger kernel size, allowing to propagate information over longer distances. Furthermore, we include experiments with the PMNet proposed in [pmnet], featuring a relatively deep encoder consisting of stacked ResNet-layers and several parallel convolutional layers with varying dilation [dilated] after the encoder. PMNet has been shown to perform better than RadioUNet and other architectures on different datasets [pmnet], [challenge]. These were the only architectures designed for the radio map prediction task available in the literature, for which the code has been made publicly available.
@@
Figure 3: Illustration (inspired by [dcn]) of sampling points in standard, dilated and deformable convolution with kernel size $3\times 3$.
@@
Beyond these state-of-the-art models, we also consider the use of deformable convolutions. Originally presented in [dcn], this CNN layer has been used in several computer vision problems, but, to the best of our knowledge, we are the first ones to apply it for radio map prediction. Similar to a dilated convolution, it allows to enlarge the receptive field by sampling the input at positions further away. The sampling points are not fixed (Fig. 3), instead, the offset compared to a standard convolution is computed from the input with learnable parameters. Intuitively, this should make it easier to propagate information in arbitrary directions compared to dilated convolutions. We use this layer as a replacement for some of the convolutions in an otherwise fairly standard UNet and, following [dcn], call the resulting model UNetDCN.
@@
Table IV compares the complexity of the different architectures across various aspects. RadioUNet is the most lightweight and the fastest model. PMNet exhibits the highest complexity in terms of multiply-accumulate operations (MACs) and number of parameters, whereas UNetDCN occupies an intermediate position between these two models.
Model RadioUNet [radiounet] PMNet [pmnet] UNetDCN
# parametersa $10.9$M $33.4$M $17,7$M
# MACsa $7.7$G $50.7$G $12.6$G
@@
TABLE IV: Complexity of the considered architectures for the default inputs and batch size 1
acalculated using DeepSpeed [deepspeed].
IV-B Input Features
@@
Following the related literature (Section II-B), we aim to encode all relevant parameters that change between the different simulations in 2D images, with each pixel representing a specific location on the map. In the following, we provide an explanation of the input features we consider. Several examples for a sample from the dataset are shown in Fig. 4, and further details of the implementation can be found in the code. All inputs are normalized to values in $[-1,1]$ before being fed to the CNN.
@@
Figure 4: Example radio map 4(a) and (normalized) input features. 4(b) Tx location (visibility enhanced, in reality only marked by one pixel), 4(c) aerial image, 4(d) nDSM, 4(e) 2D distance to Tx, 4(f) azimuth angle, 4(g), 4(h) map coordinates from GA [radiotrans], 4(i) Tx antenna pattern projected onto the ground.
@@
In the $2D$ setting, the positions of the Tx, buildings, and potentially other objects are often given in a one-hot encoding in separate binary tensors, where a $1$ indicates the presence of the Tx or object and a 00 absence in the location corresponding to each pixel (see e.g. [radiounet]). Since in our work the height of the Tx, buildings, and vegetation are relevant, we assign these as the values to the pixels instead, as it is done in [radiotrans] or [fadenet] for example (Fig. 44(b)-4(h)). The sparsity of the tensor representing the Tx location is potentially problematic, as the standard layers used in CNNs inherently have a very limited field of view, and it therefore takes several convolutional and downsampling operations to spread the information to other parts of the map.
@@
The authors of [radiotrans] propose to tackle this issue by making the information about the spatial position of each pixel together with the location of the Tx explicitly available to the model in the form of constant input tensors for the $x,y$ and $z$ coordinates of the Tx and two tensors showing the $x$ and $y$ coordinates of each pixel. Although they state that their main intention was to find an alternative to the usual positional embedding in vision transformer layers, we have found that this approach also improves the performance of other CNN models. They denote their idea as Grid Anchor (GA). Additionally, we provide the azimuth angle between the direction the Tx is pointing at and the straight line to each point of the map, as in [plnet]. This encodes Tx directivity and helps determine whether two objects lie along the same path. Lastly, we also include the distance in the $x-y$ plane from the Tx to each point as in [qiu22]. All these inputs related to positions on the maps are together denoted as coords in Sec. V.
@@
To link the antenna pattern to the spatial positions, we use spherical coordinates centered in the Tx location and rotated according to the Tx orientation to look up the gain in dB corresponding to azimuth and tilt angle for each point on the ground, as it is done in [plnet] as well.
@@
V RESULTS ON RADIO MAP PREDICTION
Model RadioUNet [radiounet] PMNet [pmnet] UNetDCN
@@ V RESULTS ON RADIO MAP PREDICTION
Image 0.094 0.0038 0.091 0.0036 0.087 0.0033
Image w/o IR $0.097$ $0.0040$ $0.093$ $0.0037$ 0.088 0.0033
Image + coords $0.098$ $0.0041$ $0.092$ $0.0036$ 0.089 0.0034
Image + nDSM $0.076$ $0.0025$ 0.0700.070{0.070}0.070 0.00210.0021{0.0021}0.0021 0.069 $\underline{\textbf{0.0020}}$
Image + nDSM + coords $\underline{0.073}$ $\underline{0.0023}$ ${\underline{0.068}}$ ${\underline{0.0020}}$ $\underline{\textbf{0.067}}$ $\underline{\textbf{0.0019}}$
TABLE V: Prediction from images – lowest errors in each column (input) without nDSM in italics, with nDSM underlined and per row (model) in bold.
@@
All models are trained with respect to mean-squared error (MSE) between the predicted and ground truth radio maps. We measure the final performance in terms of root-mean-square error (RMSE) in grayscale. Recall that the grayscale radio map values in $[0,1]$ stem from an affine transformation of the possible path loss values in dB, which lie in $[-127\text{dB},-50\text{dB}]$, hence, the RMSE in dB can be obtained from multiplying the RMSE in grayscale by a factor of $77$. Furthermore, we report the normalized mean-squared error (NMSE) after conversion to dB scale, which emphasizes the importance of samples with a high signal power (see [radiounet]).
@@
0.034
0.012
0.025
0.008
@@
0.016
0.009
0.017
0.008
0.023
0.024
0.019
0.014
0.019
0.014
(d) 0.049
0.014
(e) 0.034
0.010
(f) 0.029
0.010
Figure 5: Examples of predictions – 5(a) overlayed height maps with buildings in red, vegetation in green and Tx position in blue, 5(b) aerial image, 5(c) ground truth radio map, predictions only from images and from images, nDSMs and coords by: 5(d) RadioUNet [radiounet], 5(e) PMNet [pmnet], 5(f) UNetDCN, MSE below.
In Table V, we list the results. Overall, all models are able to estimate the radio maps with very good accuracy between $5.2$ and $7.2$ dB RMSE. Removing the infrared channel from the input image causes a slight decrease in accuracy. The height map improves the predictions significantly for all models, as expected, since inferring the height of objects from a two-dimensional image is inherently ill-posed [im2height]. Across all models, the additional inputs related to the positions on the map provide a slight improvement when also the height map is used. However, they deteriorate performance when only the image is provided, which is rather surprising. A possible explanation could be that the height information, map coordinates, and distances share the same physical units, while the channels of the aerial images contain a completely different type of information. We observe that, for all inputs considered, the more lightweight RadioUNet performs slightly worse than the two other models. The UNetDCN proposed by us achieves in all cases at least slightly better accuracy than the state-of-the-art model PMNet.
@@
Inferring the correct height of objects and their precise locations from just an image does not always work well. In the first row of Fig. 6, for example, we can see that the network without access to height information does not recognize that the house in the main lobe completely blocks the signal. Below in the second row, it seems to misinterpret the exact shape of the buildings due to shadows and predicts a non-existent propagation path.
@@
Figure 6: Examples of prediction, where height and positions of buildings are not determined correctly without nDSM. 6(a) aerial image and Tx position in blue 6(b) antenna gain projected onto the floor 6(c) ground truth radio map, predictions by PMNet [pmnet] with inputs: 6(d) only image 6(e) image and unclassified nDSM, MSE below.
VI APPLICATION TO COVERAGE OPTIMIZATION
To showcase an application of the proposed models, we consider the problem of optimizing the directivity of cellular BS to maximize the coverage in a given area. We study two possible formulations of the problem that model different scenarios, utilizing the fast interference time and differentiability of the CNN models.
VI-A Coverage Definition and Optimization Objective

### v2

@@ II-B2 Radio Map Simulations
For the Tx, we have selected the directional antenna type in the ray-tracing software. It allows to simulate an idealized main beam without any side lobes. By adjusting the parameters half power beam width and first null beam width, it can model very narrow beams or antennas covering wider sectors. For each Tx position, we test several combinations of azimuth and tilt angles in the direction pointing away from the building, on which the Tx is placed, with a line of sight algorithm. Orientations that do not cover at least a certain area on the ground are disregarded. For all angle combinations found, we select one of the narrow and one of the wide antenna patterns in Table II. Although on each city map we consider different Tx positions and orientations, each of them is used in a separate simulation, resulting in a different radio map with a single Tx per data sample.
@@
The simulation output are the magnitudes of the channel coefficients at each Rx on the grid, which we use to calculate the path loss (1). The path loss values are converted to $\operatorname{dB}$-scale and rescaled and cut following [12]. Assuming a bandwidth of $B=10\operatorname{MHz}$, thermal noise power spectral density $(N_{0})_{\operatorname{dBm}/\operatorname{Hz}}=-174\operatorname{dBm}/\operatorname{Hz}$ and an idealistic noise figure of $(\text{NF})_{\operatorname{dB}}=0\operatorname{dB}$ at the Rx (Table III), the noise floor $\mathcal{N}$ is calculated as
@@ II-B3 Other Data
We also include images from [35] taken about $2$ months after the LiDAR measurements, cut and downscaled to match the position and resolution of the nDSMs. Furthermore, the dataset contains 2D polygons with height attributes extracted from the nDSMs, which have been used for the ray-tracing simulations. Files containing line-of-sight information for each radio map are also included but not considered in this work.
II-C Experiment Design
II-C1 CNN-Architectures
@@
As a lightweight baseline model, we use the RadioUNet [12], more concretely the first part of the WNet described by the authors. In principle, it follows the structure of the original UNet [12], but it features more down and upsampling layers and, in some parts, convolutions with a larger kernel size, allowing to propagate information over longer distances. As a second baseline, we include experiments with the PMNet proposed in [5], featuring a relatively deep encoder consisting of stacked ResNet-layers and several parallel convolutional layers with varying dilation [17] after the encoder. PMNet has been shown to perform better than RadioUNet and other architectures on different datasets [5], [36]. These were the only architectures designed for the radio map prediction task available in the literature, for which the code has been made publicly available.
@@
Figure 3: Illustration (inspired by [25]) of sampling points in standard, dilated and deformable convolution with kernel size $3\times 3$.
@@
We propose the use of a model dubbed UNetDCN that combines the structure of UNet [13] with deformable convolutions. Originally presented in [25], this CNN layer has been used in several computer vision problems, but, to the best of our knowledge, we are the first ones to apply it for radio map prediction. Similar to a dilated convolution, it allows to enlarge the receptive field by sampling the input at positions further away. The sampling points are not fixed (Fig. 3), instead, the offset compared to a standard convolution is computed from the input with learnable parameters. Intuitively, this should make it easier to propagate information in arbitrary directions compared to dilated convolutions. Furthermore, in comparison to standard convolutions with a large kernel size, deformable convolutions require less parameters and numeric operations. The architecture is illustrated in Fig. 4 and the exact implementation can be found on our GitHub page mentioned in Section I. Many aspects of the network are similar to the original UNet [13], such as repeated convolutions increasing the channel dimension and max pooling reducing the spatial dimension in the encoder part, upsampling and simultaneous reduction of the channel dimension in the decoder part and skip connections from the encoder to the decoder at several stages to preserve information. The main changes on the other hand are the replacement of some standard convolutions by deformable convolutions, additional batch normalization and residual connections. In Section III-A, we provide an ablation study investigating the choice of the hyperparameters in Fig. 4 and the use of deformable convolutions.
[…]
Figure 4: Outline of the UNetDCN architecture. The numbers on the left describe the spatial dimensions and the numbers on top the channel dimension of the tensors, represented as gray blocks. In our experiments, $H=W=256$, $C_{out}=1$, $C_{in}$ depends on the chosen inputs and the width and depth hyperparameters are set to $C=32,d=3$. Blocks of network layers are depicted by arrows, and explained below. The channel dimension is always changed by the first convolution inside of each block. The layers used are standard 3$\times$3 convolutions (Conv), batch normalization (BN), rectified linear unit (ReLU), residual connections (res), concatenation of tensors along the channel dimension (cat), downsampling with $2\times 2$ MaxPool and upsampling with $2\times 2$ transposed convolutions (ConvT).
@@
Table IV compares the complexity of the different architectures across various aspects. PMNet exhibits the highest complexity in terms of multiply-accumulate operations (MACs) and number of parameters. The proposed UNetDCN has the smallest number of parameters, resulting in a minimal memory footprint, while its MAC count is slightly higher than that of RadioUNet.
Model RadioUNet [12] PMNet [5] UNetDCN
#Paramsa 10.9M 33.4M 4.5M
#MACsa 7.7G 50.7G 10.3G
@@
TABLE IV: Complexity of the considered architectures for the default inputs and batch size 1.aCalculated using DeepSpeed [37].
II-C2 Input Features
@@
Following the related literature (Section I-A), we aim to encode all relevant parameters that change between the different simulations in 2D images, with each pixel representing a specific location on the map. In the following, we provide an explanation of the input features we consider. Several examples for a sample from the dataset are shown in Fig. 5, and further details of the implementation can be found in the code. All inputs are normalized to values in $[-1,1]$ before being fed to the CNN.
@@
Figure 5: Example radio map 5(a) and (normalized) input features. 5(b) Tx location (visibility enhanced, in reality only marked by one pixel), 5(c) aerial image, 5(d) nDSM, 5(e) 2D distance to Tx, 5(f) azimuth angle, 5(g), 5(h) map coordinates from GA [16], 5(i) Tx antenna pattern projected onto the ground.
@@
In the $2D$ setting, the positions of the Tx, buildings, and potentially other objects are often given in a one-hot encoding in separate binary tensors, where a $1$ indicates the presence of the Tx or object and a 00 absence in the location corresponding to each pixel (see e.g. [12]). Since in our work the height of the Tx, buildings, and vegetation are relevant, we assign these as the values to the pixels instead, as it is done in [16] or [15] for example (Fig. 55(b)-5(h)). The sparsity of the tensor representing the Tx location is potentially problematic, as the standard layers used in CNNs inherently have a very limited field of view, and it therefore takes several convolutional and downsampling operations to spread the information to other parts of the map.
@@
The authors of [16] propose to tackle this issue by making the information about the spatial position of each pixel together with the location of the Tx explicitly available to the model in the form of constant input tensors for the $x,y$ and $z$ coordinates of the Tx and two tensors showing the $x$ and $y$ coordinates of each pixel. Although they state that their main intention was to find an alternative to the usual positional embedding in vision transformer layers, we have found that this approach also improves the performance of other CNN models. They denote their idea as Grid Anchor (GA). Additionally, we provide the azimuth angle between the direction the Tx is pointing at and the straight line to each point of the map, as in [14]. This encodes Tx directivity and helps determine whether two objects lie along the same path. Lastly, we also include the distance in the $x-y$ plane from the Tx to each point as in [21]. All these inputs related to positions on the maps are together denoted as coords in Sec. III.
@@
To link the antenna pattern to the spatial positions, we use spherical coordinates centered in the Tx location and rotated according to the Tx orientation to look up the gain in dB corresponding to azimuth and tilt angle for each point on the ground, as it is done in [14] as well.
@@
III RESULTS ON RADIO MAP PREDICTION
Model RadioUNet [12] PMNet [5] UNetDCN
@@ III RESULTS ON RADIO MAP PREDICTION
Image $\textit{0.094}(0.073/0.093)$ 0.0038 $\textit{0.091}(0.068/0.089)$ 0.0036 0.088 (0.077/0.086) 0.0033
Image w/o IR $0.097(0.070/0.094)$ $0.0040$ $0.093(0.078/0.092)$ $0.0037$ 0.089 (0.074/0.085) 0.0034
Image + coords $0.098(0.076/0.096)$ $0.0041$ $0.092(0.070/0.090)$ $0.0036$ 0.089 (0.072/0.085) 0.0034
Image + nDSM $0.076(0.062/0.071)$ $0.0025$ ${0.070}(0.057/0.066)$ 0.00210.0021{0.0021}0.0021 0.069 (0.061/0.065) 0.0020
Image + nDSM + coords $\underline{0.073}(0.063/0.070)$ $\underline{0.0023}$ ${\underline{0.068}}(0.060/0.065)$ ${\underline{0.0020}}$ 0.067 (0.061/0.064) 0.0019
TABLE V: Errors on the test set (in brackets on the training/validation set). Lowest test errors in each column (input) without nDSM in italics, with nDSM underlined and per row (model) in bold.
@@
All models are trained with respect to mean-squared error (MSE) between the predicted and ground truth radio maps. We measure the final performance in terms of root-mean-square error (RMSE) in grayscale. Recall that the grayscale radio map values in $[0,1]$ stem from an affine transformation of the possible path loss values in dB, which lie in $[-127\text{dB},-50\text{dB}]$, hence, the RMSE in dB can be obtained from multiplying the RMSE in grayscale by a factor of $77$. Furthermore, we report the normalized mean-squared error (NMSE) after conversion to dB scale, which emphasizes the importance of samples with a high signal power (see [12]).
@@
Figure 6: Errors per epoch on the training and on the validation set. Dots mark epochs in which the learning rate is reduced due to stagnating validation loss. The final model weights stem from the epoch with the lowest validation loss, which is marked with a cross.
@@
0.006
(d) 0.009
0.005
(e) 0.019
0.004
(f) 0.006
0.003
Figure 7: Examples of predictions – 7(a) overlayed height maps with buildings in red, vegetation in green and Tx position in blue, 7(b) aerial image, 7(c) ground truth radio map, predictions only from images and from images, nDSMs and coords by: 7(d) RadioUNet [12], 7(e) PMNet [5], 7(f) UNetDCN, MSE below.
In Table V, we list the results. Overall, all models are able to estimate the radio maps with very good accuracy between $5.2$ and $7.2$ dB RMSE. Removing the infrared channel from the input image causes a slight decrease in accuracy. The height map improves the predictions significantly for all models, as expected, since inferring the height of objects from a two-dimensional image is inherently ill-posed [38]. Across all models, the additional inputs related to the positions on the map provide a slight improvement when also the height map is used. However, they deteriorate performance when only the image is provided, which is rather surprising. A possible explanation could be that the height information, map coordinates, and distances share the same physical units, while the channels of the aerial images contain a completely different type of information. We observe that, for all inputs considered, the more lightweight RadioUNet performs slightly worse than the two other models. The UNetDCN we propose achieves comparable accuracy when height information is included and outperforms the baselines when height data is absent.
@@
In Fig. 6, we plot the averaged training and validation losses for each epoch in the two most relevant scenarios. We observe that without height information, the gap between training and validation loss increases stronger, which implies overfitting. The reason for this could be that inferring the correct height of objects and their precise locations from just an image is very hard and the images do not contain sufficient information for stronger generalization. UNetDCN demonstrates better resistance to overfitting, as shown by the smaller gap between training and validation/test loss. In the first row of Fig. 8, for example, we can see that the network without access to height information does not recognize that the house in the main lobe completely blocks the signal. Below in the second row, it seems to misinterpret the exact shape of the buildings due to shadows and predicts a non-existent propagation path.
@@
Figure 8: Examples of prediction, where height and positions of buildings are not determined correctly without nDSM. 8(a) aerial image and Tx position in blue 8(b) antenna gain projected onto the floor 8(c) ground truth radio map, predictions by PMNet [5] with inputs: 8(d) only image 8(e) image and unclassified nDSM, MSE below.
III-A Ablation Study
We have performed additional experiments to investigate the effect of the complexity of the proposed UNetDCN in Fig. 4 on the achieved accuracy and to evaluate the usage of deformable convolutions (DC). For the two most relevant input features, Table VI lists the training and validation error together with the complexity in terms of the number of parameters and MACs for the baseline model and modified versions.
Reducing the model complexity, either by decreasing depth ($d$) or width (number of channels, $C$), increases the validation loss notably by up to $0.004$. On the other hand, increasing the number of channels ($C$) only slightly improves validation loss at the cost of a significant increase in computational demand. The gap between training and validation loss is much larger when height information is absent. This could potentially mean a higher risk of overfitting. For deeper networks ($d=3,4$), we observe slight improvements when predicting from images alone. However, when height information is included, validation loss remains the same or slightly worsens. Overall, the proposed parameters $d=3$ and $C=32$ provide a good validation accuracy at a low complexity. Increasing the size of the model could lead to marginal improvements but at the cost of significantly higher computational complexity and, in some cases, potentially a larger risk of overfitting.
To assess the impact of deformable convolutions (DC), we also experiment with models following the same structure illustrated in Fig. 4, but with all DC layers replaced by other modules. We consider two alternatives: First, standard convolutions with larger kernel sizes and second, the atrous spatial pyramid pooling (ASPP) block from [5]. The ASPP block is an inception module with parallel $3\times 3$ convolutions at different dilation rates (Fig. 3). With standard convolutions, even with large kernel sizes it is not possible to reach the same accuracy as with DC, despite a high complexity. Increasing the kernel size beyond $9\times 9$ provides no further benefit but rather leads to performance deterioration. Also the ASPP block does not work well as a drop-in replacement in our proposed architecture. As we observe signs of potential overfitting in the first column of Table VI, we also try two more variants with reduced complexity in terms of the parameters $d$ and $C$, but the performance is still subpar compared to DC. These findings support our hypothesis that DC are a very efficient way to increase the receptive field in the radio map prediction task.
RMSE
Modification Image +nDSM #Params #MACs
+coords
- 0.077/0.086 0.061/0.064 4.5M 10.3G
$C=16$ 0.082/0.090 0.066/0.067 1.19 M 3.15G
$C=64$ 0.069/0.085 0.059/0.063 17.56M 36.9G
$d=2$ 0.081/0.088 0.067/0.068 1.15M 8.02G
$d=4$ 0.073/0.085 0.062/0.064 17.7M 12.6G
$d=5$ 0.072/0.084 0.063/0.065 69.8M 14.7G
Conv $5\times 5$ 0.083/0.090 0.067/0.070 7.05M 31.3G
Conv $7\times 7$ 0.078/0.089 0.067/0.069 11.2M 53.5G
Conv $9\times 9$ 0.073/0.088 0.064/0.067 16.7M 83.0G
Conv $11\times 1$ 11 0.074/0.088 0.067/0.069 23.7M 120G
Conv $13\times 13$ 0.074/0.090 0.065/0.068 32.0M 164G
ASPP [5] 0.072/0.093 0.063/0.069 8.62M 38.7G
ASPP, $C=16$ 0.088/0.097 0.070/0.073 2.2M 9.72G
ASPP, $d=2$ 0.074/0.091 0.067/0.070 2.1M 27.2G
TABLE VI: Comparison of training/validation loss and complexity in terms of the number of parameters and MACs for different variations of UNetDCN. Results for the proposed model with no modification are listed in the first row for comparison.
IV APPLICATION TO COVERAGE OPTIMIZATION
To showcase an application of the proposed models, we consider the problem of optimizing the directivity of cellular BS to maximize the coverage in a given area. We study two possible formulations of the problem that model different scenarios, utilizing the fast interference time and differentiability of the CNN models. More precisely, we freeze the network after training and treat it as a differentiable function from the inputs, that include a representation of the antenna orientation which we aim to optimize, to the generated radio maps. Via backpropagation through the frozen network and gradient descent, we can optimize the antenna orientation parameters with respect to the coverage, that is defined as a function of the radio maps.
IV-A Coverage Definition and Optimization Objective